# Shape Formation and Locomotion with Joint Movements in the Amoebot Model

### v2

@@ 3.3 Stability
One might argue that removing too many bonds may destabilize the structure during the movements. It is therefore desirable to maintain the connectivity between the amoebots as much as possible. Ideally, we want to only remove a constant fraction of all incident bonds of a connected component. Formally, let $\Gamma_{G}(U)=\{\{u,v\}\in E\mid u\in U\land v\not\in U\}$ where $G=(V,E)$. For each connected component $U\subseteq S$, we require $|\Gamma_{G_{R}}(U)|=\Theta(|\Gamma_{G_{S}}(U)|)$. This already holds for all presented movement primitives except for the slide and $k$-tunnel primitives of rhombical meta-modules. In the following, we present alternative implementations that increase the connectivity for these primitives.
First, consider the slide primitive. Recall that we need to slide a meta-module $R_{1}$ along two meta-modules $R_{2}$ and $R_{3}$ (see Figure 5(c)). We refer to Section 3.1 for details. The idea is to enlarge the moving part as follows (see Figure 12). For the sake of simplicity, we assume that $\ell$ is divisible by $8$. We partition $R_{2}$ and $R_{3}$ into $4\times 4$ submodules, respectively. The slide primitive consists of the following three rounds. In the first round, we utilize the submodules adjacent to the start and target position of $R_{1}$ to move $R_{1}$ to the center position. In the second round, we restore $R_{2}$ and $R_{3}$ again. In the last two rounds, we mirror the first two rounds to move $R_{1}$ to its target position.
@@
Our alternative implementation of the slide primitive requires $3$ rounds.
Second, consider the $k$-tunnel primitive. Recall that we need to move a meta-module $R$ through a simple path of meta-modules with $k$ corners to the other end (see Figure 5(e)). We refer to Section 3.1 for details.
@@
In the first round, we contract or expand all submodules that we would have in the original version of the $k$-tunnel primitive – except for the ones at the two ends of the line. We use those two to maintain the bonds to attached meta-modules. Note that if we maintained the bonds at other submodules, we would move the line.
@@
3.4 Meta-Module Construction
First, we form a line of $r$ rhombical meta-modules of side length $\ell-1$ from a line of $n=r\cdot\ell^{2}/2$ amoebots as follows (see Figure 14(a)). For that, we divide the line of amoebots into groups of $\ell$ amoebots and apply the following procedure introduced by Feldmann et al. [23] on each group of amoebots. The amoebots expand such that the height is doubled. Next, we reorientate the amoebots using handovers. Then, the amoebots contract again such that the width is halved. Thus, each iteration of the procedure doubles the height and halves the width. We repeat this procedure until the width is 1. Clearly, we need $O(\log\ell)$ iterations. Each iteration requires $O(1)$ rounds. Overall, the procedure requires $O(\log\ell)$ rounds. At the end of the second step, we obtain a parallelogram of side length $\ell-1$ and $r\cdot\ell/2-1$ filled with contracted amoebots. It remains to expand all amoebots to obtain the line of $r$ rhombical meta-modules. We obtain the following theorem.
Theorem 3.12.
There is a centralized algorithm that forms a line of $r$ rhombical meta-modules of side length $\ell-1$ from a line of $n=r\cdot\ell^{2}/2$ amoebots within $O(\log n)$ rounds.
(a) Rhombical meta-modules.
(b) Hexagonal meta-modules.
Figure 15: Construction of meta-modules. Adjacent amoebots and meta-modules of the same color belong to the same group, respectively.
Second, we form a line of $h=\frac{r}{3}$ hexagonal meta-modules of alternating side lengths $\ell$ and $\ell-1$ from a line of $r$ rhombical meta-modules of side length $\ell-1$ as follows (see Figure 14(b)). At the start, we divide the line of rhombical meta-modules into groups of $3$ rhombical meta-modules and apply the following three steps to each group. In the first step, we apply the $k$-tunnel primitive to move the first meta-module through the second meta-module. In the second step, we apply the realignment primitive on the first and third meta-module. In the third step, we apply the procedure from Section 4.1 to align the first and third meta-module with the hexagonal tesselation. Alternatively, we could also apply a modified slide primitive in order to keep the expansion as high as possible. The modification consists in the fact that we only contract a single amoebot such that the meta-module is only moved by a distance of 1. The procedure requires $O(1)$ rounds since all steps require $O(1)$ rounds. We obtain the following theorem.
Theorem 3.13.
There is a centralized algorithm that forms a line of $h=\frac{r}{3}$ hexagonal meta-modules of alternating side lengths $\ell$ and $\ell-1$ from a line of $r$ rhombical meta-modules of side length $\ell-1$ within $O(\log n)$ rounds.
4 Shape Formation
In this section, we discuss possible shape formation algorithms. For that, we look at shape formation algorithms for other lattice-type MRSs. In particular, we consider the polylogarithmic time solutions for the nubot model [60] and the crystalline atom model [6]. The first subsection deals with the former model while the second subsection deals with shape formation algorithms for our meta-modules including the one for latter model.
@@ 4.1 Nubot Model
Similar to the amoebot model, the nubot model [60] considers robots on the triangular grid where at most one robot can be positioned on each node. Adjacent robots can be connected by rigid bonds [^2] Woods et al. [60] distinguish between rigid and flexible bonds. We ignore that since the flexible bonds are not necessary to achieve the polylogarithmic time shape formation algorithm.. In the joint movement extension, the rigid bonds correspond to $E_{R}$.
@@
For each connected component $C$, we perform the following steps in parallel. If there is an amoebot $u\in C$ that is adjacent to the same amoebot $v\in M$ before and after the translation, we proceed as follows (see Figure 15(a)). Let $A$ denote the row of amoebots in $C$ through $u$ into direction $d^{\prime\prime}$. In the first move cycle, we remove all bonds between $C$ and $M$ except for the bond between $u$ and $v$, and each amoebot in $A$ expands into direction $d^{\prime}$. In the second move cycle, we remove all bonds between $C$ and $M$ except for the new bond between $u$ and $v$, and each amoebot in $A$ contracts. Note that both movements are possible due to 1.
If there is no amoebot in $C$ that is adjacent to the same amoebot in $M$ before and after the translation, we proceed as follows (see Figure 15(b)). Let $B$ denote all amoebots in $C$ that have an unoccupied node in direction $d^{\prime}$. In the first move cycle, each amoebot in $B$ expands into direction $d^{\prime}$. There has to be an amoebot $u\in B$ that becomes adjacent to an amoebot $v\in M$. Otherwise, the resulting amoebot structure would not be connected. In the second move cycle, we remove all bonds between $C$ and $M$ except for the bond between $u$ and $v$, and each amoebot in $B$ contracts. Note that both movements are possible due to 1.
@@
Woods et al. [60] showed that in the nubot model, the robots are able to self-assemble arbitrary shapes/patterns in an amount of time equal to the worst-case running time for a Turing machine to compute a pixel in the shape/pattern plus an additional factor which is polylogarithmic in its size. While we are able to perform the translations, we do not have the means to let amoebots appear and disappear in the amoebot model. This prevents us from simulating the shape formation algorithm by Woods et al. [60].
4.2 Shape Formation Algorithms for Meta-Modules
Naturally, our meta-modules allow us to simulate shape formation algorithms for lattice-type MRSs of similar shape if we can implement the same movement primitives. This leads us to the following results.
@@ 4.2 Shape Formation Algorithms for Meta-Modules
There is a centralized shape formation algorithm for $m$ rhombical meta-modules that requires $O(\log m)$ rounds and performs $\Theta(m\log m)$ moves overall.
@@
Aloupis et al. [6] proposed a shape formation algorithm for crystalline atoms. It requires $O(\log m)$ rounds and performs $\Theta(m\log m)$ moves overall. The idea is to transform the initial shape to a canonical shape using a divide and conquer approach. The target shape is reached by reversing that procedure. We refer to [6] for the details. The algorithm utilizes the contraction, slide and tunnel primitives which our rhombical meta-modules are capable of (see Section 3.1). Hence, they can simulate this shape formation algorithm.
@@
There is a centralized shape formation algorithm for $m$ hexagonal meta-modules that requires $O(m)$ rounds. Each module has to perform at most $O(m)$ moves.
@@
Hurtado et al. [27] proposed a shape formation algorithm for hexagonal robots. It requires $O(m)$ rounds and each module has to perform at most $O(m)$ moves. The idea is to compute a spanning tree and to move the robots along the boundary of the tree to a leader module where the robots form a canonical shape, e.g., a line. The target shape is reached by reversing that procedure. We refer to [27] for the details. The algorithm utilizes the rotation primitive which our hexagonal meta-modules are capable of (see Section 3.2). Hence, they can simulate this shape formation algorithm.
@@ 5 Locomotion
In this section, we consider amoebot structures capable of locomotion along an even surface. There are three basic types of terrestrial locomotion: rolling, crawling, and walking [26, 33]. We can find biological and artificial examples for each of those. In the following subsections, we will present an amoebot structure for each type and analyze their velocity. In the last subsection, we discuss the transportation of objects.
@@ 5.2 Crawling
The simplest amoebot structure that imitates an earthworm is a simple line of $n$ expanded amoebots along the surface (see Figure 17(a)). Each amoebot can be seen as a segment of the worm. Instead of propagating waves of contractions and expansions, we contract and expand the whole structure at once. During each contraction (expansion), we release all bonds between the amoebot structure and the surface except for the ones at the head (tail) of the structure that serve as an anchor. As a result, the contraction (expansion) pulls (pushes) the structure to the front. The simple line has moved by a distance of $n$ along the surface after performing a contraction and an expansion. This is the fastest way possible to move along a surface since we accumulate the movements of all amoebots into the same direction. By performing the movements periodically, we obtain the following theorem.
@@
(a) Simple line of amoebots.
(b) Line of rhombical meta-modules.
(c) Millipede structure. The red rhombi indicate the legs. The blue and green rhombi indicate the body. The blue rhombi indicate the connectors. Note that the figure only depicts the first half of a period.
Figure 18: Worm and millipede structures.
However, in practice, the contractions and expansions of the whole structure yield high forces acting on the connections within the amoebot structure. We can address this problem by thickening the worm structure. This increases the expansion of the structure and with that its stability. Consider a line of $r$ rhombical meta-modules of side length $\ell-1$ (see Figure 17(b)). Each module can be seen as a segment of the worm structure. Recall that we can contract a rhombical meta-module into a parallelogram (see Figure 7(a)). The line of rhombical meta-modules moves in the same manner as the simple line except for the following two points. First, we utilize the whole meta-module at the front and at the end as an anchor to increase the grip, respectively. Second, only the middle $r-2$ meta-modules participate in the contractions and expansions. The line of rhombical meta-modules moves a distance of $\frac{r-2}{2}\cdot\ell$ along the surface after performing a contraction and an expansion. By performing the movements periodically, we obtain the following theorem.
@@
Another problem in practice is friction between the structure and the surface and with that the wear of the structure. The worm structure is therefore poorly scalable in its length such that other types for locomotion are more suitable for large amounts of amoebots.
@@ 5.3 Walking
Our walking amoebot structure imitates millipedes. Millipedes have flexible, segmented bodies with tens to hundreds of legs that provide morphological robustness [52]. They move by propagating leg-density waves from the posterior to the anterior [25, 37]. We build our millipede structure from rhombical meta-modules of side length $\ell-1$ (see Figure 17(c)). Let $p$ denote the number of legs. The body and each leg consists of a line of rhombical meta-modules. All legs have the same size. Let $q$ denote the number of rhombical modules in each leg. We attach each leg to a meta-module of the body and orientate them alternating to the front and to the back. We call those meta-modules connectors. In order to prevent the legs from colliding, we place $q$ meta-modules between two connectors. Altogether, the millipede structure consists of $(2p-1)\cdot q+p$ meta-modules.
In order to move the legs back and forth, we simply apply the realignment primitive (see Figure 7(c)) on all meta-modules within the legs (see Figure 17(c)). We achieve forward motion by releasing all bonds between the surface and the legs moving forwards. In one step, the body moves a distance of $q\cdot\ell$. Note that the number of legs has no impact on the velocity. By continuously repeating these leg movements, we achiegfve a motion similar to the leg-density waves of millipedes. Note that we reach the initial amoebot structure after two leg movements. Hence, we obtain the following theorem.
@@
Our millipede structure composed of rhombical meta-modules of side length $\ell-1$ with $p$ legs composed of $q$ rhombical meta-modules moves a distance of $2\cdot q\cdot\ell$ within each period of constant length.

### v3

@@ 3.4 Stability
One might argue that removing too many bonds may destabilize the structure during the movements. It is therefore desirable to maintain the connectivity between the amoebots as much as possible. In the following, we propose two measures for stability.
First, we consider the connectivity of the amoebot structure during each movement. Note that we only decrease the connectivity in the first step of each movement where we may remove bonds. In the second and fourth step, the connectivity does not change, and in the third step, we can only increase the connectivity by adding new bonds. Ideally, we want to just remove a constant fraction of all incident bonds of a connected component.
Formally, let $\Gamma_{G}(U)=\{\{u,v\}\in E\mid u\in U\land v\not\in U\}$ where $G=(V,E)$. Recall that $S_{i}$ denotes the amoebot structure before the first step, and $S_{i}^{0}$ the amoebot structure $S_{i}$ after the first step.
Definition 3.17.
We call a reconfiguration $(S_{0},\dots,S_{F})$ locally stable if and only if
$\forall i\in\{0,\dots,F-1\}:|\Gamma_{S_{i}^{0}}(U)|=\Theta(|\Gamma_{S_{i}}(U)|)\text{ for all connected components }U\subseteq S_{i}.$
Remark 3.18.
There are locally stable reconfigurations $(S_{0},\dots,S_{F})$ such that the reverse reconfiguration $(S_{F},\dots,S_{0})$ is not locally stable. For example, consider the second step of the second last phase of our algorithm that forms a line of $r$ rhombical meta-modules of side length $\ell-1$ from a line of $n=r\cdot\ell^{2}/2$ amoebots (see Theorems 3.10 and 21). For $r=1$, we have $\ell=O(\sqrt{n})$. Recall that each group consists of $\ell$ amoebots. The reconfiguration only removes a constant fraction of the bonds between two groups and adds $O(\sqrt{n})$ bonds afterwards. Hence, the reverse reconfiguration has to remove these $O(\sqrt{n})$ bonds except for one.
Second, we consider the connectivity of the amoebot structure during the whole reconfiguration. Ideally, a stable reconfiguration algorithm should keep the connectivity as high as possible during the whole process. The weakest point of a reconfiguration $(S_{0},\dots,S_{F})$ is the amoebot structure $S_{i}$ with the lowest connectivity. However, the achievable connectivity highly depends on the initial amoebot structure $S_{0}$ and the target amoebot structure $S_{F}$. Hence, the connectivity of $S_{i}$ should be at least as high as the one of $S_{0}$ and $S_{F}$ with the lower connectivity. Recall that $\Gamma_{G}(U)=\{\{u,v\}\in E\mid u\in U\land v\not\in U\}$ where $G=(V,E)$. Let
$\alpha(G)=\min_{U\subseteq V,1\leq|U|\leq|V|/2}\frac{|\Gamma(U)|}{|U|}.$
Definition 3.19.
We call a reconfiguration $(S_{0},\dots,S_{F})$ globally stable if and only if
$\min_{i\in\{0,\dots,F\}}\alpha(G_{S_{i}})=\Omega(\min\{\alpha(G_{S_{0}}),\alpha(G_{S_{F}})\}).$
Definition 3.20.
We call a reconfiguration $(S_{0},\dots,S_{F})$ stable if and only if it is locally and globally stable.
Note that both constructions in Section 3.3 are stable. In order to obtain stable reconfigurations in general, locally stable movement primitives can be helpful. All so far presented movement primitives are locally stable except for the slide, rotation and $k$-tunnel primitives for rhombical meta-modules, and the switching primitive for hexagonal meta-modules. In the following, we present alternative implementations that increase the connectivity for these primitives.
First, consider the slide primitive. Recall that we need to slide a meta-module $R_{1}$ along two meta-modules $R_{2}$ and $R_{3}$ (see Figure 8(c)). We refer to Section 3.1 for details. The idea is to enlarge the moving part as follows (see Figure 23). For the sake of simplicity, we assume that $\ell$ is divisible by $8$. We partition $R_{2}$ and $R_{3}$ into $4\times 4$ submodules, respectively. The slide primitive consists of the following four rounds. In the first round, we utilize the submodules adjacent to the start and target position of $R_{1}$ to move $R_{1}$ to the center position. In the second round, we restore $R_{2}$ and $R_{3}$ again. In the last two rounds, we mirror the first two rounds to move $R_{1}$ to its target position.
@@
Our alternative implementation of the slide primitive requires $4$ rounds.
Second, consider the rotation and $k$-tunnel primitives. Since the rotation primitive is a special case of the $k$-tunnel primitive, we will constrain the explanation to the latter. Recall that we need to move a meta-module $R$ through a simple path of meta-modules with $k$ corners to the other end (see Figure 8(e)). We refer to Section 3.1 for details.
@@
In the first round, we contract or expand all submodules that we would have in the original version of the $k$-tunnel primitive – except for the ones at the two ends of the line. We use those two to maintain the bonds to attached meta-modules.
@@
Corollary 3.23.
Our alternative implementation of the rotation primitive requires $6$ rounds.
Remark 3.24.
We can utilize our alternative implementation of the rotation primitive to make the algorithms that form a line of hexagonal meta-modules locally stable (see Theorems 3.13 and 3.15). Note that the first step of the rotation primitive, i.e., the contraction of the meta-modules, is already locally stable (see first round in Figure 22). Hence, it suffices to only use the alternative implementation on the second step, i.e., the expansion of the meta-modules (see second round in Figure 22). Therefore, the required number of rounds of Theorems 3.13 and 3.15 increases by $2$ to $10$ and $22$, respectively.
Finally, consider the switching primitive. Recall that we need to switch between the two possibilities to arrange the rhombical meta-modules in a hexagonal meta-module (see Figure 18). We refer to Section 3.2 for details.
Consider our implementation in Figure 18. Note that only the second round does not satisfy our connectivity requirement. Hence, we alter the way we exchange the smaller triangles between the rhombical meta-modules as follows. We replace the second and third round in Figure 18 with the four rounds in Figure 26. The idea is to apply a similar technique as in the reorientation primitive for rhombical meta-modules. Instead of exchanging all amoebots at once, we exchange them in two phases. In the first phase, we exchange each second pair of amoebots on each line between the bigger triangles (see first and second round in Figure 26). In the second phase, we exchange all other pairs of amoebots (see third and fourth round in Figure 26). Note that we can again combine the first two (first round in Figure 18 and first round in Figure 26) and last two rounds (last round in Figure 18 and last round in Figure 26), respectively.
Figure 26: Switching primitive with higher expansion.
Lemma 3.25.
Our alternative implementation of the switching primitive requires $4$ rounds.
4 Reconfiguration
In this section, we discuss possible reconfiguration algorithms. For that, we look at reconfiguration algorithms for other lattice-type MRSs and discuss whether amoebots are capable of simulating these. In particular, we consider the polylogarithmic time solutions for the nubot model [62] and the crystalline atom model [6]. The first subsection deals with the former model while the second subsection deals with reconfiguration algorithms for our meta-modules including the one for the latter model. None of the reconfiguration algorithms discussed in this section will be stable.
@@ 4.1 Nubot Model
Similar to the amoebot model, the nubot model [62] considers robots on the triangular grid where at most one robot can be positioned on each node. Adjacent robots can be connected by rigid bonds [^2] Woods et al. [62] distinguish between rigid and flexible bonds. We ignore that since the flexible bonds are not necessary to achieve the polylogarithmic time reconfiguration algorithm.. In the joint movement extension, the rigid bonds correspond to the edge set $B$.
@@
For each connected component $C$, we perform the following steps in parallel. If there is an amoebot $u\in C$ that is adjacent to the same amoebot $v\in M$ before and after the translation, we proceed as follows (see Figure 27(a)). Let $A$ denote the row of amoebots in $C$ through $u$ into direction $d^{\prime\prime}$. Note that $A$ may only contain $u$. In the first round, we remove all bonds between $C$ and $M$ except for the bond between $u$ and $v$, and each amoebot in $A$ expands into direction $d^{\prime}$. In the second round, we remove all bonds between $C$ and $M$ except for the new bond between $u$ and $v$, and each amoebot in $A$ contracts. Note that both movements are possible due to 1.
If there is no amoebot in $C$ that is adjacent to the same amoebot in $M$ before and after the translation, we proceed as follows (see Figure 27(b)). Let $B$ denote all amoebots in $C$ that have an unoccupied node in direction $d^{\prime}$. In the first round, each amoebot in $B$ expands into direction $d^{\prime}$. There has to be an amoebot $u\in B$ that becomes adjacent to an amoebot $v\in M$. Otherwise, the resulting amoebot structure would not be connected. In the second round, we remove all bonds between $C$ and $M$ except for the bond between $u$ and $v$, and each amoebot in $B$ contracts. Note that both movements are possible due to 1.
@@
Woods et al. [62] showed that in the nubot model, the robots are able to self-assemble arbitrary shapes/patterns in an amount of time equal to the worst-case running time for a Turing machine to compute a pixel in the shape/pattern plus an additional factor which is polylogarithmic in its size. While we are able to perform the translations, we do not have the means to let amoebots appear and disappear in the amoebot model. This prevents us from simulating the reconfiguration algorithm by Woods et al. [62].
4.2 Reconfiguration Algorithms for Meta-Modules
Naturally, our meta-modules allow us to simulate reconfiguration algorithms for lattice-type MRSs of similar shape if we can implement the same movement primitives. This leads us to the following results.
@@ 4.2 Reconfiguration Algorithms for Meta-Modules
There is a centralized reconfiguration algorithm for $m$ rhombical meta-modules that requires $O(\log m)$ rounds and performs $\Theta(m\log m)$ moves overall.
@@
Aloupis et al. [6] proposed a reconfiguration algorithm for crystalline atoms. It requires $O(\log m)$ rounds and performs $\Theta(m\log m)$ moves overall. The idea is to transform the initial structure to a canonical structure using a divide and conquer approach. The target structure is reached by reversing that procedure. We refer to [6] for the details. The algorithm utilizes the contraction, slide and tunnel primitives which our rhombical meta-modules are capable of (see Section 3.1). Hence, they can simulate this reconfiguration algorithm.
@@
There is a centralized reconfiguration algorithm for $m$ hexagonal meta-modules that requires $O(m)$ rounds. Each module has to perform at most $O(m)$ moves.
@@
Hurtado et al. [28] proposed a reconfiguration algorithm for hexagonal robots. It requires $O(m)$ rounds and each module has to perform at most $O(m)$ moves. The idea is to compute a spanning tree and to move the robots along the boundary of the tree to a leader module where the robots form a canonical structure, e.g., a line. The target structure is reached by reversing that procedure. We refer to [28] for the details. The algorithm utilizes the rotation primitive which our hexagonal meta-modules are capable of (see Section 3.2). Hence, they can simulate this reconfiguration algorithm.
@@ 5 Locomotion
In this section, we consider amoebot structures capable of locomotion along an even surface. There are three basic types of terrestrial locomotion: rolling, crawling, and walking [27, 34]. We can find biological and artificial examples for each of those. In the following subsections, we will present an amoebot structure for each type and analyze their velocity. In the last subsection, we discuss the transportation of objects. All constructions are locally and globally stable unless stated otherwise.
@@ 5.2 Crawling
The simplest amoebot structure that imitates an earthworm is a simple line of $n$ expanded amoebots along the surface (see Figure 29). Each amoebot can be seen as a segment of the worm. Instead of propagating waves of contractions and expansions, we contract and expand the whole structure at once. During each contraction (expansion), we release all bonds between the amoebot structure and the surface except for the ones at the head (tail) of the structure that serve as an anchor. As a result, the contraction (expansion) pulls (pushes) the structure to the front. The simple line has moved by a distance of $n$ along the surface after performing a contraction and an expansion. This is the fastest way possible to move along a surface since we accumulate the movements of all amoebots into the same direction. By performing the movements periodically, we obtain the following theorem.
Figure 29: Simple line of amoebots.
@@
However, in practice, the contractions and expansions of the whole structure yield high forces acting on the connections within the amoebot structure. We can address this problem by thickening the worm structure. This increases the expansion of the structure and with that its stability. Consider a line of $r$ rhombical meta-modules of side length $\ell-1$ (see Figure 30). Each module can be seen as a segment of the worm structure. Recall that we can contract a rhombical meta-module into a parallelogram (see Figure 10). The line of rhombical meta-modules moves in the same manner as the simple line except for the following two points. First, we utilize the whole meta-module at the front and at the end as an anchor to increase the grip, respectively. Second, only the middle $r-2$ meta-modules participate in the contractions and expansions. The line of rhombical meta-modules moves a distance of $\frac{r-2}{2}\cdot\ell$ along the surface after performing a contraction and an expansion. By performing the movements periodically, we obtain the following theorem.
Figure 30: Line of rhombical meta-modules.
@@
Note that unless the number of amoebots resp. meta-modules is constant, neither amoebot structure in the subsection is locally stable since we have to remove almost all bonds between the structure and the surface. Another problem in practice that comes with the local instability is friction between the structure and the surface and with that the wear of the structure. The worm structure is therefore poorly scalable in its length such that other types for locomotion are more suitable for large amounts of amoebots.
@@ 5.3 Walking
Our walking amoebot structure imitates millipedes. Millipedes have flexible, segmented bodies with tens to hundreds of legs that provide morphological robustness [54]. They move by propagating leg-density waves from the posterior to the anterior [25, 38]. We build our millipede structure from rhombical meta-modules of side length $\ell-1$ (see Figure 32). Let $p$ denote the number of legs. The body and each leg consists of a line of rhombical meta-modules. All legs have the same size. Let $q$ denote the number of rhombical modules in each leg. We attach each leg to a meta-module of the body and orientate them alternating to the front and to the back. We call those meta-modules connectors. In order to prevent the legs from colliding, we place $q$ meta-modules between two connectors. Altogether, the millipede structure consists of $(2p-1)\cdot q+p$ meta-modules.
Figure 32: Millipede structure. The arrow indicates two rounds. The red rhombi indicate the legs. The blue and green rhombi indicate the body. The blue rhombi indicate the connectors. Note that the figure only depicts the first half of a period.
In order to move the legs back and forth, we simply apply the realignment primitive (see Figure 12) on all meta-modules within the legs (see Figure 32). We achieve forward motion by releasing all bonds between the surface and the legs moving forwards. In one step, the body moves a distance of $q\cdot\ell$. Note that the number of legs has no impact on the velocity. By continuously repeating these leg movements, we achieve a motion similar to the leg-density waves of millipedes. Note that we reach the initial amoebot structure after two leg movements. Hence, we obtain the following theorem.
@@
Our millipede structure composed of rhombical meta-modules of side length $\ell-1$ with $p$ legs composed of $q$ rhombical meta-modules moves a distance of $2\cdot q\cdot\ell$ every $4$ rounds.